OpenStudy (anonymous):

How do you simplify 12^log sub 12 base 144?

OpenStudy (anonymous):

\[12^{\log _{12}144}\]

jimthompson5910 (jim_thompson5910):

Hint: \[\Large b^{\log_b(x)} = x\]

OpenStudy (anonymous):

so the answer is 144?

OpenStudy (anonymous):

what about for \[9^{\log _{3}20}\]

OpenStudy (anonymous):

i know u can get it to \[3^{2^{\log _{3}20}}\]

OpenStudy (anonymous):

but then what?

jimthompson5910 (jim_thompson5910):

yes the answer to the first one is 144

OpenStudy (anonymous):

do you know about the second

jimthompson5910 (jim_thompson5910):

\[\Large 9^{\log_{3}(20)}\] \[\Large (3^2)^{\log_{3}(20)}\] \[\Large 3^{2\log_{3}(20)}\] \[\Large 3^{\log_{3}(20^2)}\]

OpenStudy (anonymous):

so once you get the base of 3 you kind of do a reverse power property? so the answer is 400?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks!!!